OpenStudy (anonymous):

2m-nx=x+4, solve for x

OpenStudy (anonymous):

@Matt80M 2m - nx = x + 4 -nx + 2m + nx = x + 4 + nx 2m = nx + x + 4 2m/2 = nx + x + 4/2 m = 1/2nx + 1/2x + 2